OpenStudy (anonymous):

solve for x. x - 7 = 7 ---- -- 8 2

OpenStudy (anonymous):

is that (x-7)/2 = 7/2?

OpenStudy (anonymous):

or (x-7)/8 = 7/2

OpenStudy (anonymous):

x-7/8 = 7/2

OpenStudy (anonymous):

i don't mind helping...but your gonna have to write it out for me properly

OpenStudy (anonymous):